# A Dynamical Theory of Overfitting and Generalization in Large Two-Layer Networks

### v2

@@ D.2 Accuracy of the numerical integration scheme
where ${\boldsymbol{P}}_{{\mathbb{S}}^{d-1}}$ is the projector to the unit sphere, i.e. ${\boldsymbol{P}}_{{\mathbb{S}}^{d-1}}({\boldsymbol{x}})={\boldsymbol{x}}/\|{\boldsymbol{x}}\|$ if ${\boldsymbol{x}}\neq{\mathbf{0}}$ and ${\boldsymbol{P}}_{{\mathbb{S}}^{d-1}}({\mathbf{0}})={\mathbf{0}}$. Note that the trajectories of Eq. (D.9) depend on the sample size $n$ (and hence the dimension $d=d_{n}$) and the stepsize $\eta_{\mbox{\tiny\rm GD}}$. We denote them by ${\boldsymbol{a}}^{n}(t;\eta_{\mbox{\tiny\rm GD}})$ $\boldsymbol{W}^{n}(t;\eta_{\mbox{\tiny\rm GD}})$.
@@ F.1.3 The algorithmic interpolation transition
Figure 10: Training with pure noise data under lazy initialization: algorithmic interpolation threshold. We plot the time for GF to converge to near-zero training error as a function of $\alpha$, for various values of $\gamma_{0}$, as computed using the theory of Section F.1.2. The divergence of $t_{\text{rel}}$ signals the phase transition for GF interpolation $\alpha_{\mbox{\tiny\rm GF}}^{*}$. Inset: $\tau_{\text{rel}}$ versus $\alpha$ in linear scale. Main panel: $t_{\text{rel}}$ versus $\alpha_{\mbox{\tiny\rm GF}}^{*}-\alpha$ (with the fitted value of $\alpha^{*}_{\mbox{\tiny\rm GF}}$). Here we use $h(z)=(3/10)z+z^{2}/2$.
@@ H.2 Infinite width limit at fixed $\overline{\alpha}$
Appendix I Lower bounding the overfitting timescale
@@
I.1 Proof of Theorem 3.1
By computing the derivative $\partial_{a_{i}}\widehat{\mathscrsfs{R}}({\boldsymbol{a}}(t),\boldsymbol{W}(t))$, we get
@@
$\sqrt{2\widehat{\mathscrsfs{R}}_{n}({\boldsymbol{a}}(0),\boldsymbol{W}(0))}$ $=C(\tau+a_{0}L)\,.$
Summarizing the above bounds
@@
${\mathcal{W}}_{m,d}({\overline{a}})$ $:=\Big{\{}({\boldsymbol{a}},\boldsymbol{W})\in{\mathbb{R}}^{m}\times{\mathbb{R}}^{m\times d}:\;\frac{\|{\boldsymbol{a}}\|_{1}}{m}\leq{\overline{a}},\;\|{\boldsymbol{w}}_{i}\|_{2}=1\;\;\forall i\leq m\Big{\}}\,$, (I.1)
${\mathcal{W}}_{m,d}^{\infty}({\overline{a}})$ $:=\Big{\{}({\boldsymbol{a}},\boldsymbol{W})\in{\mathbb{R}}^{m}\times{\mathbb{R}}^{m\times d}:\;\|{\boldsymbol{a}}\|_{\infty}\leq{\overline{a}},\;\|{\boldsymbol{w}}_{i}\|_{2}=1\;\;\forall i\leq m\Big{\}}\,.$ (I.2)
@@
$\sup_{({\boldsymbol{a}},\boldsymbol{W})\in{\mathcal{W}}_{m,d}({\overline{a}})}\big{|}\widehat{\mathscrsfs{R}}_{n}({\boldsymbol{a}},\boldsymbol{W})-\mathscrsfs{R}({\boldsymbol{a}},\boldsymbol{W})\big{|}\leq CL^{2}{\overline{a}}^{2}\sqrt{\frac{d}{n}}\,.$ (I.3)
This is a well-known result, that we reproduce for the readers’ convenience. By usual concentration, it is sufficient to bound the expectation. This in turns is bounded by symmetrization and contraction inequalities. Letting $({\varepsilon}_{i})_{i\leq n}\sim_{iid}{\sf Unif}(\{+1,-1\})$, we have
${\mathbb{E}}$ $\sup_{({\boldsymbol{a}},\boldsymbol{W})\in{\mathcal{W}}_{m,d}({\overline{a}})}\big{|}\widehat{\mathscrsfs{R}}_{n}({\boldsymbol{a}},\boldsymbol{W})-\mathscrsfs{R}({\boldsymbol{a}},\boldsymbol{W})\big{|}\leq{\mathbb{E}}\sup_{({\boldsymbol{a}},\boldsymbol{W})\in{\mathcal{W}}_{m,d}({\overline{a}})}\frac{1}{n}\sum_{i=1}^{n}{\varepsilon}_{i}\big{(}y_{i}-f({\boldsymbol{x}}_{i};{\boldsymbol{a}},\boldsymbol{W})\big{)}^{2}$
$\leq 2{\mathbb{E}}\sup_{({\boldsymbol{a}},\boldsymbol{W})\in{\mathcal{W}}_{m,d}({\overline{a}})}\frac{1}{n}\sum_{i=1}^{n}{\varepsilon}_{i}y_{i}f({\boldsymbol{x}}_{i};{\boldsymbol{a}},\boldsymbol{W})+{\mathbb{E}}\sup_{({\boldsymbol{a}},\boldsymbol{W})\in{\mathcal{W}}_{m,d}({\overline{a}})}\frac{1}{n}\sum_{i=1}^{n}{\varepsilon}_{i}f({\boldsymbol{x}}_{i};{\boldsymbol{a}},\boldsymbol{W})^{2}$
@@
$E_{1}=$ ${\mathbb{E}}\sup_{({\boldsymbol{a}},\boldsymbol{W})\in{\mathcal{W}}_{m,d}({\overline{a}})}\sum_{j=1}^{m}\frac{a_{j}}{m}\frac{1}{n}\sum_{i=1}^{n}{\varepsilon}_{i}y_{i}\sigma({\boldsymbol{w}}_{j}^{{\sf T}}{\boldsymbol{x}}_{i})$
$\leq$ ${\overline{a}}\,{\mathbb{E}}\sup_{\|{\boldsymbol{w}}\|=1}\frac{1}{n}\sum_{i=1}^{n}{\varepsilon}_{i}y_{i}\sigma({\boldsymbol{w}}^{{\sf T}}{\boldsymbol{x}}_{i})$
$\stackrel{{\scriptstyle(a)}}{{\leq}}$ ${\overline{a}}\,L{\mathbb{E}}\sup_{\|{\boldsymbol{w}}\|=1}\frac{1}{n}\sum_{i=1}^{n}{\varepsilon}_{i}(1+|y_{i}|){\boldsymbol{w}}^{{\sf T}}{\boldsymbol{x}}_{i}$
$\leq$ ${\overline{a}}\,L{\mathbb{E}}\Big{\{}\Big{\|}\frac{1}{n}\sum_{i=1}^{n}{\varepsilon}_{i}(1+|y_{i}|){\boldsymbol{x}}_{i}\Big{\|}\Big{\}}$
$\leq$ $C{\overline{a}}\,L\sqrt{\frac{d}{n}}\,$,
@@
$E_{2}=$ ${\mathbb{E}}\sup_{({\boldsymbol{a}},\boldsymbol{W})\in{\mathcal{W}}_{m,d}({\overline{a}})}\sum_{j,l=1}^{m}\frac{a_{j}a_{l}}{m^{2}}\frac{1}{n}\sum_{i=1}^{n}{\varepsilon}_{i}\sigma({\boldsymbol{w}}_{j}^{{\sf T}}{\boldsymbol{x}}_{i})\sigma({\boldsymbol{w}}_{l}^{{\sf T}}{\boldsymbol{x}}_{i})$
$\leq{\overline{a}}^{2}{\mathbb{E}}\sup_{{\boldsymbol{w}},\tilde{\boldsymbol{w}}\in{\mathbb{S}}^{d-1}}\frac{1}{n}\sum_{i=1}^{n}{\varepsilon}_{i}\sigma({\boldsymbol{w}}^{{\sf T}}{\boldsymbol{x}}_{i})\sigma(\tilde{\boldsymbol{w}}^{{\sf T}}{\boldsymbol{x}}_{i})$
$\stackrel{{\scriptstyle(b)}}{{\leq}}CL^{2}{\overline{a}}^{2}{\mathbb{E}}\sup_{{\boldsymbol{w}}\in{\mathbb{S}}^{d-1}}\frac{1}{n}\sum_{i=1}^{n}{\varepsilon}_{i}{\boldsymbol{w}}^{{\sf T}}{\boldsymbol{x}}_{i}$
@@
${\mathbb{E}}$ $\sup_{({\boldsymbol{a}},\boldsymbol{W})\in{\mathcal{W}}_{m,d}({\overline{a}})}\big{|}\widehat{\mathscrsfs{R}}_{n}({\boldsymbol{a}},\boldsymbol{W})-\mathscrsfs{R}({\boldsymbol{a}},\boldsymbol{W})\big{|}\leq CL^{2}{\overline{a}}^{2}\sqrt{\frac{d}{n}}\,$,
and Eq. (I.3) follows by a concentration argument.
I.2 Proof of Theorem 3.2
@@
${\boldsymbol{g}}^{{\boldsymbol{w}}}_{n,\ell}({\boldsymbol{a}},\boldsymbol{W})$ $:=\frac{m}{|a_{\ell}|}\big{[}\nabla_{{\boldsymbol{w}}_{\ell}}\widehat{\mathscrsfs{R}}_{n}({\boldsymbol{a}},\boldsymbol{W})-\nabla_{{\boldsymbol{w}}_{\ell}}\mathscrsfs{R}({\boldsymbol{a}},\boldsymbol{W})\big{]}\,$, (I.4)
$g^{a}_{n,\ell}({\boldsymbol{a}},\boldsymbol{W})$ $:=m\big{[}\nabla_{a_{\ell}}\widehat{\mathscrsfs{R}}_{n}({\boldsymbol{a}},\boldsymbol{W})-\nabla_{a_{\ell}}\mathscrsfs{R}({\boldsymbol{a}},\boldsymbol{W})\big{]}\,.$ (I.5)
@@
$\sup_{({\boldsymbol{a}},\boldsymbol{W})\in{\mathcal{W}}_{m,d}({\overline{a}})}\max_{\ell\leq m}\big{\|}{\boldsymbol{g}}^{{\boldsymbol{w}}}_{n,\ell}({\boldsymbol{a}},\boldsymbol{W})\big{\|}\leq C(L^{2}{\overline{a}}+\tau^{2})\sqrt{\frac{d}{n}\log(ne/d)}\,$, (I.6)
$\sup_{({\boldsymbol{a}},\boldsymbol{W})\in{\mathcal{W}}_{m,d}({\overline{a}})}\max_{\ell\leq m}\big{|}g^{a}_{n,\ell}({\boldsymbol{a}},\boldsymbol{W})\big{|}\leq C(L^{2}{\overline{a}}+\tau^{2})\sqrt{\frac{d}{n}}\,.$ (I.7)
@@
$\leq 2\,{\mathbb{E}}\sup_{{\boldsymbol{w}},{\boldsymbol{u}}}\frac{1}{n}\sum_{i=1}^{n}{\varepsilon}_{i}y_{i}\sigma^{\prime}({\boldsymbol{w}}^{{\sf T}}{\boldsymbol{x}}_{i}){\boldsymbol{u}}^{{\sf T}}{\boldsymbol{x}}_{i}+2\,{\overline{a}}{\mathbb{E}}\sup_{{\boldsymbol{w}},\overline{\boldsymbol{w}},{\boldsymbol{u}}}\frac{1}{n}\sum_{i=1}^{n}{\varepsilon}_{i}\sigma^{\prime}({\boldsymbol{w}}^{{\sf T}}{\boldsymbol{x}}_{i})\sigma(\overline{\boldsymbol{w}}^{{\sf T}}{\boldsymbol{x}}_{i}){\boldsymbol{u}}^{{\sf T}}{\boldsymbol{x}}_{i}$

### v3

@@ D.1 Integration technique
The discretization of Eq. (C.10) deserves an additional clarification because of the delta-function. For $t_{a}\geq t_{b}$, $t_{a},t_{b}\in{\mathbb{N}}\eta$, we compute
$\frac{R_{ij}(t_{a}+\eta,t_{b})-R_{ij}(t_{a},t_{b})}{\eta}=-\nu_{i}(t_{a})R_{ij}(t_{a},t_{b})-\frac{1}{\eta}\delta_{ij}{\boldsymbol{1}}_{t_{a}=t_{b}}$
$-\frac{1}{m}\sum_{l=1}^{m}\sum_{s\in[t_{a},t_{b}]\cap{\mathbb{N}}\eta}M^{R}_{il}(t_{a},s)\,R_{lj}(s,t_{b})\,\eta\,$,
with boundary condition
$R_{ij}(t_{b},t_{b})$ $=0\,\;\;\;\;\forall i,j\leq m\,.$
@@ D.2 Accuracy of the numerical integration scheme
where ${\boldsymbol{P}}_{{\mathbb{S}}^{d-1}}$ is the projector to the unit sphere, i.e. ${\boldsymbol{P}}_{{\mathbb{S}}^{d-1}}({\boldsymbol{x}})={\boldsymbol{x}}/\|{\boldsymbol{x}}\|$ if ${\boldsymbol{x}}\neq{\mathbf{0}}$ and ${\boldsymbol{P}}_{{\mathbb{S}}^{d-1}}({\mathbf{0}})={\mathbf{0}}$. Note that the trajectories of Eq. (D.9) depend on the sample size $n$ (and hence the dimension $d=d_{n}$) and the stepsize $\eta_{\mbox{\tiny\rm GD}}$. To emphasize this dependence, we also use the notation ${\boldsymbol{a}}^{n}(t;\eta_{\mbox{\tiny\rm GD}})$ $\boldsymbol{W}^{n}(t;\eta_{\mbox{\tiny\rm GD}})$.
@@ F.1.3 The algorithmic interpolation transition
Figure 10: Training with pure noise data under lazy initialization: algorithmic interpolation threshold. Left Panel. We plot the train error as a function of time for different values of $\alpha$ at $\gamma_{0}=10/7$. The train error has an exponential decay to zero for $\alpha$ below the interpolation threshold. Right Panel. We plot the time for GF to converge to near-zero training error as a function of $\alpha$, for various values of $\gamma_{0}$, as computed using the theory of Section F.1.2. The divergence of $t_{\text{rel}}$ signals the phase transition for GF interpolation $\alpha_{\mbox{\tiny\rm GF}}^{*}$. Inset: $\tau_{\text{rel}}$ versus $\alpha$ in linear scale. Main panel: $t_{\text{rel}}$ versus $\alpha_{\mbox{\tiny\rm GF}}^{*}-\alpha$ (with the fitted value of $\alpha^{*}_{\mbox{\tiny\rm GF}}$). Here we use $h(z)=(3/10)z+z^{2}/2$.
@@ H.2 Infinite width limit at fixed $\overline{\alpha}$
Appendix I Details about SGD simulations
In this appendix we provide some details about the numerical simulations with stochastic gradient descent (SGD) presented in Figures 2, 4.
We generate data according to the pure noise model $y_{i}={\varepsilon}_{i}$ (Fig. 2), $y_{i}=\varphi({\boldsymbol{w}}_{*}^{{\sf T}}{\boldsymbol{x}}_{i})+{\varepsilon}_{i}$ (Fig. 4), $i\leq n$. We learn the two-layer network of Eq. (1.1), see below for the class definition.
class Net($n$.Module):
def __init__(self, a, m, d):
super().__init__()
self.m = m
self.lin1 = $n$.Linear(d,m,bias=False)
self.lin1.weight.data = (1/np.sqrt(d))*torch.randn((m,d))
self.lin2 = $n$.Linear(m,1,bias=False)
self.lin2.weight.data[0,:] = a
self.act = Myact()
self.project()
def forward(self, x ):
x1 = self.act(self.lin1(x))
return self.lin2(x1)/self.m
def project(self, epsilon):
row_norms = torch.norm(self.lin1.weight.data, dim=1, keepdim=True)
row_norms = torch.clamp(row_norms, min=epsilon)
self.lin1.weight.data = self.lin1.weight.data/row_norms
As shown in this code, we use the initialization
$\big({\boldsymbol{a}}_{0},\boldsymbol{W}_{0}\big)$ $={\boldsymbol{P}}_{{\sf B}}\big(\overline{{\boldsymbol{a}}}_{0},\overline{\boldsymbol{W}}_{0}\big)\,$, (I.1)
$\overline{{\boldsymbol{a}}}_{0}$ $=(a_{0},\dots,a_{0})\,,\;\;\;(W_{0,ij})_{i\leq m,j\leq d}\sim{\sf N}(0,1/d)\,.$ (I.2)
where ${\boldsymbol{P}}_{{\sf B}}$ projects first layer weights to the unit ball:
${\boldsymbol{P}}_{{\sf B}}\Big({\boldsymbol{a}},({\boldsymbol{w}}_{1},\dots,{\boldsymbol{w}}_{m})\Big)=\left({\boldsymbol{a}},\Big(\frac{{\boldsymbol{w}}_{1}}{\|{\boldsymbol{w}}_{1}\|\wedge 1},\dots,\frac{{\boldsymbol{w}}_{m}}{\|{\boldsymbol{w}}_{m}\|\wedge 1}\Big)\right)\,.$ (I.3)
We use the standard SGD iteration without weight decay and constant stepsize $\eta$, and batch size $b$:
$\overline{{\boldsymbol{\theta}}}_{k+1}$ $={\boldsymbol{\theta}}_{k}-\eta\nabla\widehat{\mathscrsfs{R}}_{S(k)}({\boldsymbol{\theta}}_{k})\,,\;\;\;\;\widehat{\mathscrsfs{R}}_{S}({\boldsymbol{\theta}})=\frac{1}{2|S|}\sum_{i\in S}\big(y_{i}-f({\boldsymbol{x}}_{i};{\boldsymbol{\theta}})\big)^{2}\,$, (I.4)
${\boldsymbol{\theta}}_{k+1}$ $={\boldsymbol{P}}_{{\sf B}}(\overline{\boldsymbol{\theta}}_{k+1})\,.$ (I.5)
The optimizer is defined in the code below
optimizer = optim.SGD(net.parameters(), lr=lr, momentum=0., weight_decay=0.)
lambda_step = lambda epoch: 1
scheduler = torch.optim.lr_scheduler.LambdaLR(optimizer, lr_lambda=lambda_step)
In the simulations of Figures 2, and 4 we use batch size $b=100$ and step size $\eta=0.1$. Each symbol reports the average of $N_{\text{sim}}=10$ simulations.
Appendix J Lower bounding the overfitting timescale
@@ Appendix I Details about SGD simulations
J.1 Proof of Theorem 3.1
By computing the derivative $\partial_{a_{i}}\widehat{\mathscrsfs{R}}_{n}({\boldsymbol{a}}(t),\boldsymbol{W}(t))$, we get
@@
$\sqrt{2\widehat{\mathscrsfs{R}}_{n}({\boldsymbol{a}}(0),\boldsymbol{W}(0))}$ $=C(\tau+\sqrt{k}+a_{0}L)\,.$
Summarizing the above bounds, we have
@@
${\mathcal{W}}_{m,d}({\overline{a}})$ $:=\Big\{({\boldsymbol{a}},\boldsymbol{W})\in{\mathbb{R}}^{m}\times{\mathbb{R}}^{m\times d}:\;\frac{\|{\boldsymbol{a}}\|_{1}}{m}\leq{\overline{a}},\;\|{\boldsymbol{w}}_{i}\|_{2}=1\;\;\forall i\leq m\Big\}\,$, (J.1)
${\mathcal{W}}_{m,d}^{\infty}({\overline{a}})$ $:=\Big\{({\boldsymbol{a}},\boldsymbol{W})\in{\mathbb{R}}^{m}\times{\mathbb{R}}^{m\times d}:\;\|{\boldsymbol{a}}\|_{\infty}\leq{\overline{a}},\;\|{\boldsymbol{w}}_{i}\|_{2}=1\;\;\forall i\leq m\Big\}\,.$ (J.2)
@@
$\sup_{({\boldsymbol{a}},\boldsymbol{W})\in{\mathcal{W}}_{m,d}({\overline{a}})}\big|\widehat{\mathscrsfs{R}}_{n}({\boldsymbol{a}},\boldsymbol{W})-\mathscrsfs{R}({\boldsymbol{a}},\boldsymbol{W})\big|\leq C(L^{2}{\overline{a}}^{2}+\tau^{2})\sqrt{\frac{d}{n}}\,.$ (J.3)
This is a standard estimate, that we reproduce for the readers’ convenience.
We begin by bounding the expectation of the supremum by symmetrization and contraction inequalities. Letting $(\xi_{i})_{i\leq n}\sim_{iid}{\sf Unif}(\{+1,-1\})$, we have
${\mathbb{E}}$ $\sup_{({\boldsymbol{a}},\boldsymbol{W})\in{\mathcal{W}}_{m,d}({\overline{a}})}\big|\widehat{\mathscrsfs{R}}_{n}({\boldsymbol{a}},\boldsymbol{W})-\mathscrsfs{R}({\boldsymbol{a}},\boldsymbol{W})\big|\leq{\mathbb{E}}\sup_{({\boldsymbol{a}},\boldsymbol{W})\in{\mathcal{W}}_{m,d}({\overline{a}})}\frac{1}{n}\sum_{i=1}^{n}\xi_{i}\big(y_{i}-f({\boldsymbol{x}}_{i};{\boldsymbol{a}},\boldsymbol{W})\big)^{2}$
$\leq 2{\mathbb{E}}\sup_{({\boldsymbol{a}},\boldsymbol{W})\in{\mathcal{W}}_{m,d}({\overline{a}})}\frac{1}{n}\sum_{i=1}^{n}\xi_{i}y_{i}f({\boldsymbol{x}}_{i};{\boldsymbol{a}},\boldsymbol{W})+{\mathbb{E}}\sup_{({\boldsymbol{a}},\boldsymbol{W})\in{\mathcal{W}}_{m,d}({\overline{a}})}\frac{1}{n}\sum_{i=1}^{n}\xi_{i}f({\boldsymbol{x}}_{i};{\boldsymbol{a}},\boldsymbol{W})^{2}$
@@
$E_{1}=$ ${\mathbb{E}}\sup_{({\boldsymbol{a}},\boldsymbol{W})\in{\mathcal{W}}_{m,d}({\overline{a}})}\sum_{j=1}^{m}\frac{a_{j}}{m}\frac{1}{n}\sum_{i=1}^{n}\xi_{i}y_{i}\sigma({\boldsymbol{w}}_{j}^{{\sf T}}{\boldsymbol{x}}_{i})$
$\leq$ ${\overline{a}}\,{\mathbb{E}}\sup_{\|{\boldsymbol{w}}\|=1}\frac{1}{n}\sum_{i=1}^{n}\xi_{i}y_{i}\sigma({\boldsymbol{w}}^{{\sf T}}{\boldsymbol{x}}_{i})$
$\stackrel{{\scriptstyle(a)}}{{\leq}}$ ${\overline{a}}\,L{\mathbb{E}}\sup_{\|{\boldsymbol{w}}\|=1}\frac{1}{n}\sum_{i=1}^{n}\xi_{i}(1+|y_{i}|){\boldsymbol{w}}^{{\sf T}}{\boldsymbol{x}}_{i}$
$\leq$ ${\overline{a}}\,L{\mathbb{E}}\Big\{\Big\|\frac{1}{n}\sum_{i=1}^{n}\xi_{i}(1+|y_{i}|){\boldsymbol{x}}_{i}\Big\|\Big\}$
$\leq$ $C{\overline{a}}\,L(L+\tau)\sqrt{\frac{d}{n}}\,$,
@@
$E_{2}=$ ${\mathbb{E}}\sup_{({\boldsymbol{a}},\boldsymbol{W})\in{\mathcal{W}}_{m,d}({\overline{a}})}\sum_{j,l=1}^{m}\frac{a_{j}a_{l}}{m^{2}}\frac{1}{n}\sum_{i=1}^{n}\xi_{i}\sigma({\boldsymbol{w}}_{j}^{{\sf T}}{\boldsymbol{x}}_{i})\sigma({\boldsymbol{w}}_{l}^{{\sf T}}{\boldsymbol{x}}_{i})$
$\leq{\overline{a}}^{2}{\mathbb{E}}\sup_{{\boldsymbol{w}},\tilde{\boldsymbol{w}}\in{\mathbb{S}}^{d-1}}\frac{1}{n}\sum_{i=1}^{n}\xi_{i}\sigma({\boldsymbol{w}}^{{\sf T}}{\boldsymbol{x}}_{i})\sigma(\tilde{\boldsymbol{w}}^{{\sf T}}{\boldsymbol{x}}_{i})$
$\stackrel{{\scriptstyle(b)}}{{\leq}}CL^{2}{\overline{a}}^{2}{\mathbb{E}}\sup_{{\boldsymbol{w}}\in{\mathbb{S}}^{d-1}}\frac{1}{n}\sum_{i=1}^{n}\xi_{i}{\boldsymbol{w}}^{{\sf T}}{\boldsymbol{x}}_{i}$
@@
${\mathbb{E}}$ $\sup_{({\boldsymbol{a}},\boldsymbol{W})\in{\mathcal{W}}_{m,d}({\overline{a}})}\big|\widehat{\mathscrsfs{R}}_{n}({\boldsymbol{a}},\boldsymbol{W})-\mathscrsfs{R}({\boldsymbol{a}},\boldsymbol{W})\big|\leq C(L^{2}{\overline{a}}^{2}+\tau^{2})\sqrt{\frac{d}{n}}\,.$ (J.4)
In order to complete the proof of Eq. (J.3), we will show that the supremum concentrates around its expectation. For fixed $({\boldsymbol{a}},\boldsymbol{W})\in{\mathcal{W}}_{m,d}({\overline{a}})$, we have
$\big|f({\boldsymbol{x}};{\boldsymbol{a}},\boldsymbol{W})-f({\boldsymbol{x}}^{\prime};{\boldsymbol{a}},\boldsymbol{W})\big|$ $\leq L{\overline{a}}\|{\boldsymbol{x}}-{\boldsymbol{x}}^{\prime}\|_{2}\,$,
$\big|\varphi({\boldsymbol{U}}^{{\sf T}}{\boldsymbol{x}})-\varphi({\boldsymbol{U}}^{{\sf T}}{\boldsymbol{x}}^{\prime})\big\|$ $\leq L\|{\boldsymbol{x}}-{\boldsymbol{x}}^{\prime}\|_{2}\,.$
We write $\widehat{\mathscrsfs{R}}_{n}(\boldsymbol{X};{\boldsymbol{a}},\boldsymbol{W})$ to emphasize the dependence of the risk on $\boldsymbol{X}$ Letting $r({\boldsymbol{x}}_{i};{\boldsymbol{a}},\boldsymbol{W})=\varphi({\boldsymbol{U}}^{{\sf T}}{\boldsymbol{x}})-f({\boldsymbol{x}};{\boldsymbol{a}},\boldsymbol{W})$, we have
$\nabla_{{\boldsymbol{x}}_{i}}\widehat{\mathscrsfs{R}}_{n}(\boldsymbol{X};{\boldsymbol{a}},\boldsymbol{W})$ $=\frac{1}{n}\big({\varepsilon}_{i}+r({\boldsymbol{x}}_{i};{\boldsymbol{a}},\boldsymbol{W})\big)\nabla_{{\boldsymbol{x}}_{i}}r({\boldsymbol{x}}_{i};{\boldsymbol{a}},\boldsymbol{W})\,$,
$\Rightarrow\big\|\nabla_{{\boldsymbol{x}}_{i}}\widehat{\mathscrsfs{R}}_{n}(\boldsymbol{X};{\boldsymbol{a}},\boldsymbol{W})\big\|$ $\leq\frac{C}{n}(|{\varepsilon}_{i}|+L{\overline{a}})L{\overline{a}}\,.$
Hence
$\big\|\nabla_{\boldsymbol{X}}\widehat{\mathscrsfs{R}}_{n}(\boldsymbol{X};{\boldsymbol{a}},\boldsymbol{W})\big\|$ $\leq\frac{C}{\sqrt{n}}L{\overline{a}}\Big(L{\overline{a}}+\frac{\|{\boldsymbol{{\varepsilon}}}\|}{\sqrt{n}}\Big)$
$\leq\frac{C^{\prime}}{\sqrt{n}}L{\overline{a}}(L{\overline{a}}+\tau)\,$,
where the last inequality holds on an event that has probability at least $1-e^{-n}$. Defining $Z_{n,d,m}({\overline{a}}):=\sup_{({\boldsymbol{a}},\boldsymbol{W})\in{\mathcal{W}}_{m,d}({\overline{a}})}\big|\widehat{\mathscrsfs{R}}_{n}({\boldsymbol{a}},\boldsymbol{W})-\mathscrsfs{R}({\boldsymbol{a}},\boldsymbol{W})\big|$, Borell inequality yields
${\mathbb{P}}\Big\{\big|Z_{n,d,m}({\overline{a}})-{\mathbb{E}}Z_{n,d,m}({\overline{a}})\big|\geq B\,t\Big\}\leq 2\,e^{-nt^{2}}+e^{-n}\,$,
$B:=C^{\prime\prime}L{\overline{a}}(L{\overline{a}}+\tau)\,.$
Together with Eq. (J.4), we thus obtain that the following holds with probability $1-2e^{-t}-e^{-n}$
${\mathbb{E}}$ $\sup_{({\boldsymbol{a}},\boldsymbol{W})\in{\mathcal{W}}_{m,d}({\overline{a}})}\big|\widehat{\mathscrsfs{R}}_{n}({\boldsymbol{a}},\boldsymbol{W})-\mathscrsfs{R}({\boldsymbol{a}},\boldsymbol{W})\big|\leq C(L^{2}{\overline{a}}^{2}+\tau^{2})\sqrt{\frac{d}{n}}+C(L^{2}{\overline{a}}^{2}+\tau^{2})\sqrt{\frac{t}{n}}\,.$
This yields the desired claim.
J.2 Proof of Theorem 3.2
@@
${\boldsymbol{g}}^{{\boldsymbol{w}}}_{n,\ell}({\boldsymbol{a}},\boldsymbol{W})$ $:=\frac{m}{|a_{\ell}|}\big[\nabla_{{\boldsymbol{w}}_{\ell}}\widehat{\mathscrsfs{R}}_{n}({\boldsymbol{a}},\boldsymbol{W})-\nabla_{{\boldsymbol{w}}_{\ell}}\mathscrsfs{R}({\boldsymbol{a}},\boldsymbol{W})\big]\,$, (J.5)
$g^{a}_{n,\ell}({\boldsymbol{a}},\boldsymbol{W})$ $:=m\big[\nabla_{a_{\ell}}\widehat{\mathscrsfs{R}}_{n}({\boldsymbol{a}},\boldsymbol{W})-\nabla_{a_{\ell}}\mathscrsfs{R}({\boldsymbol{a}},\boldsymbol{W})\big]\,.$ (J.6)
@@
$\sup_{({\boldsymbol{a}},\boldsymbol{W})\in{\mathcal{W}}_{m,d}({\overline{a}})}\max_{\ell\leq m}\big\|{\boldsymbol{g}}^{{\boldsymbol{w}}}_{n,\ell}({\boldsymbol{a}},\boldsymbol{W})\big\|\leq C(L^{2}{\overline{a}}+\tau^{2})\sqrt{\frac{d}{n}\log(ne/d)}\,$, (J.7)
$\sup_{({\boldsymbol{a}},\boldsymbol{W})\in{\mathcal{W}}_{m,d}({\overline{a}})}\max_{\ell\leq m}\big|g^{a}_{n,\ell}({\boldsymbol{a}},\boldsymbol{W})\big|\leq C(L^{2}{\overline{a}}+\tau^{2})\sqrt{\frac{d}{n}}\,.$ (J.8)
@@
$\leq 2\,{\mathbb{E}}\sup_{{\boldsymbol{w}},{\boldsymbol{u}}}\frac{1}{n}\sum_{i=1}^{n}\xi_{i}y_{i}\sigma^{\prime}({\boldsymbol{w}}^{{\sf T}}{\boldsymbol{x}}_{i}){\boldsymbol{u}}^{{\sf T}}{\boldsymbol{x}}_{i}+2\,{\overline{a}}{\mathbb{E}}\sup_{{\boldsymbol{w}},\overline{\boldsymbol{w}},{\boldsymbol{u}}}\frac{1}{n}\sum_{i=1}^{n}\xi_{i}\sigma^{\prime}({\boldsymbol{w}}^{{\sf T}}{\boldsymbol{x}}_{i})\sigma(\overline{\boldsymbol{w}}^{{\sf T}}{\boldsymbol{x}}_{i}){\boldsymbol{u}}^{{\sf T}}{\boldsymbol{x}}_{i}$